# Machine Unlearning: A Comprehensive Survey

### v1

@@ 4.3. Bayesian-based Unlearning
Different from certified data removal that unlearns samples by subtracting corresponding Hessian matrix estimation from trained models, Bayesian-based unlearning tries to optimize an approximate posterior of the remaining dataset only using the trained model (the posterior of the original full dataset) and the erased data. The exact Bayesian unlearning posterior can be derivated from the Bayesian rule as $p(\theta|D_{r})=p(\theta|D)\ p(D_{e}|D_{r})/p(D_{e}|\theta)$, where $\theta$ is the posterior (i.e., model parameters). $D$ is the original full trainset, and $D_{r}$ is the remaining dataset which removes the erased data $D_{e}$, denoted as $D=D_{e}\cup D_{r}$ and $D_{e}\cap D_{r}=\emptyset$. If the model parameters $\theta$ are discrete-valued, $p(\theta|D_{r})$ can be directly obtained from the Bayesian rule. The use of a conjugate prior furthermore makes unlearning relatively simple.
@@
However, in practice, it is not easy to achieve the exact posterior, not to mention the unlearning posterior. In (Nguyen et al., 2020), Nguyen et al. tried doing likewise at the beginning. They defined the loss function using the KLD between the approximate predictive distribution $q_{u}(y|D_{r})$ and the exact predictive distribution $p(y|D_{r})$ where the observation $y$ (i.e., drawn from a model with parameters $\theta$) is conditionally independent of $D_{r}$ given $\theta$. They bounded this loss function by the KLD between posterior beliefs $q_{u}(\theta|D_{r})$ and $p(\theta|D_{r})$ and further proposed evidence upper bound (EUBO) as the loss function to unlearn the approximate unlearning posterior. To avoid the overestimation of using KL divergence to optimize the posterior, they introduced an adjusted likelihood to control the unlearning extent.
@@
In (Nguyen et al., 2022a), the authors also studied the problem of ” unlearning” particular erased subset samples from a trained model with better efficiency than retraining a new model from scratch. Toward this purpose, Nguyen et al. proposed a Markov chain Monte Cario-based machine unlearning (MCU) method deriving from the Bayesian rule. They experimentally proved that MCU could unlearn the erased subsets of the whole training dataset from a prepared model effectively and efficiently.
@@
Fu et al. (Fu et al., 2022) converted the Markov chain Monte Carlo unlearning problem into an explicit optimization problem. Then they proposed $\epsilon-$Knowledge removal, which was a little similar to certified removal methods, but they defined that the KLD between the unlearned and the retrained model must be less than $\epsilon$. To quantify the explicit $\epsilon-$knowledge removal, they proposed a knowledge removal estimator to evaluate the difference between the original and unlearned distributions. As they indicated, though their algorithm cannot wholly remove the learned knowledge from the already trained distribution, their method can still help the unlearned model approach its local minimum.
5. Distributed and Irregular Data Unlearning
We have introduced popular centralized machine unlearning in the former two sections. Most of them are requested to delete the regularly structured data. Their main purpose is to reduce the computation and storage costs and the accuracy degradation caused by unlearning. Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in distributed and irregular data learning scenarios. In this section, we will introduce the literature about federated unlearning and graph unlearning as two representative directions.
5.1. Federated Unlearning
Federated learning (McMahan et al., 2016, 2017) is first proposed to protect the participating users’ privacy during a machine learning training process in distributed situations. All participants will only upload their locally trained model parameters instead of their sensitive local data to the server for the model training. Therefore, in a federated learning scenario, the limited access to the dataset will become a unique challenge to implement unlearning.
Since the local data cannot be uploaded to the FL server side, most federated unlearning methods tried to unlearn a certain client’s influence from the trained model by storing and estimating the contribution of uploaded parameters. In this situation, they can implement federated unlearning without interacting with the client. The two representative methods are (Liu et al., 2021; Wu et al., 2022b). Liu et al. (Liu et al., 2021) proposed FedEraser to sanitize the contribution of a federated client on the global federated learning model. In particular, during the FL global training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates. Based on these operations, they reconstructed the unlearned global model instead of retraining a new model from scratch. However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data. It is unsuitable for a client who wants to unlearn a small piece of his data. Wu et al. (Wu et al., 2022b) tried to unlearn a client’s influence from the federated global model by removing the historical updates from the global model. They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
Wang et al. (Wang et al., 2022) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL. The biggest challenge is that the local data used for FL global model’s training cannot be accessed globally. Therefore, they explored the inner influence of each channel in deep and observed that various channels have distinct impacts on different categories. They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model. A method called Term Frequency Inverse Document Frequency (TF-IDF) was introduced to quantize the class discrimination of channels. Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them. Evaluations of CIFAR10 and CIFAR100 demonstrate the unlearning effect of their method.
Different from unlearning a whole client’s influence and unlearning a class, Liu et al. (Liu et al., 2022a) investigated how to unlearn data samples in FL. They first defined a federated unlearning problem and proposed a fast retraining method to remove data points from the federated learning global model. Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods (Xu et al., 2020) and the first-order Taylor approximate technique. They used the practical Fisher Information Matrix (FIM) (Ly et al., 2017) to model the Hessian matrix at a low cost. When unlearning, they subtract the estimation matrix from gradients with a bound.
5.2. Graph Unlearning
We introduce graph unlearning as a representative kind of irregular data unlearning. In (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022), researchers extend regular data machine unlearning to a graph data scenario. Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig. 6. Therefore, Chien et al. (Chien et al., 2022) proposed node feature unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC). Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees. They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
@@ 5.2. Graph Unlearning
Figure 6. (a) Regular data unlearning, which only unlearns the data sample. (b) Graph unlearning. Since the graph data contains both node features and connective edge information, we may need to unlearn only edges information or both edges and nodes in graph unlearning.
@@
Chen et al. (Chen et al., 2021a) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation. Therefore, they proposed GraphEraser to implement unlearning tailored to graph data. Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs. Then they trained the constituent graph models on these sub-graphs and ensemble them for the final prediction task. To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
@@
In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN. To unlearn the knowledge of a specific node, they propose a projection-based unlearning approach PROJECTOR that projects the weight parameters of the pre-trained model onto a subspace that is irrelevant to the deleted node features. PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
@@ 6. Unlearning Evaluation and Verification
Table 2. Evaluation and Verification Metrics, Datasets
Evaluation Metrics Formula/Description Usage Literature Popular Datasets
Accuracy Accuracy of the unlearned model on erased datasets and remaining datasets To evaluate the predictive accuracy of the unlearned model (Bourtoule et al., 2021; Guo et al., 2020), … MNIST (LeCun et al., 1998) CIFAR SVHN LSUN Photo (Shchur et al., 2018) ImageNet (Deng et al., 2009) Adult Credit info PID (Asuncion and Newman, 2007) PWD (Liu et al., 2010) AG News (Zhang et al., 2015) GTSRB (Stallkamp et al., 2011) Covtype (Blackard and Dean, 1999) HIGGS (Baldi et al., 2014) RCV1 (Glewis et al., 2004) Market-1501 (Zheng et al., 2015) YELP2018 Movielens-1m Movielens-10m Cora (Sen et al., 2008) Citseer (Yang et al., 2016) Pubmed (Fey and Lenssen, 2019) ogbn-arxiv (Hu et al., 2020) Computers (McAuley et al., 2015) CS (Shchur et al., 2018) Physics (Shchur et al., 2018)
@@
KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(A(D_{r})||U(D_{e},A(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
JS-Divergence The JS divergence between the predictions of retrained and unlearned model: $JS(A,U)=0.5\cdot KL(A||Q)+0.5\cdot KL(U||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
Membership inference Recall (##\##detected items /#absent#/\#/ #erased items) To verify if the erased data is unlearned by the model (Graves et al., 2021)
Forgetting rate (FR) $FR=\frac{AF-BF}{BT}$, where AF, BF and BT are defined below To measure the rate of samples that are changed from member to non -member after unlearning (Ma et al., 2022)
@@
MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries to the model To verify if the model unlearns the backdoored data (Hu et al., 2022)
@@
Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in unlearning. In Section 2, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics. The common metrics also include the accuracy to assess the predictive accuracy of the unlearned model and running time to evaluate the unlearning efficiency. We have listed recent evaluation and verification metrics in Table 6, where the popular experimental datasets are furthermore presented. Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here. In this section, we pay attention to the new evaluation metrics that recent studies tailored to unlearning. Most of them are based on privacy leakage.
@@
In (Liu and Tsaftaris, 2020), Liu and Tsaftaris indicated that the data auditing problem in machine unlearning is becoming popular. They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed. Their challenge was that the method would fail when the calibration dataset was out of high quality or when the query samples were similar to the training dataset. An Ensembled Membership Auditing (EMA) (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model. EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools. In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset. In the second step, EMA aggregates a final auditing score by ensembling many membership metrics and statistical tools. They verified whether a trained model forgets query data by the final auditing score. They conducted the experiments using standard datasets, including MNIST, SVHN, and the Chest X-ray dataset.
@@
Inspired by backdoor attacks in ML, Hu et al. (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB). MIB leverage the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong class. The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if a model has been trained based on her data. MIB evaluate the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model. Then, they observed the results between the targeted model and clean models to infer if the model was trained using the backdoored data.
@@
A similar membership-inference-based method was proposed in (Ma et al., 2022). Ma et al. verified the effectiveness of an unlearning method by their proposed uniform metric called forgetting rate (FR). Their methods were inspired by membership inference attacks and defined the evaluation metrics based on the observation of membership inference. Suppose the dataset $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$. In their FR definition, $BF$ and $AF$ mean the number of samples in $D_{e}$ that are predicted to be false by a membership inference attack before and after unlearning. $BT$ is the number of samples in $D_{e}$ predicted to be true by a membership inference attack before machine unlearning. According to this definition, $FR$ presents an instinctive evaluation of how many data points are changed from member to non-member after unlearning. If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective. By contrast, the unlearning will be meaningless. They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets. The main idea of Forsaken is to continuously generate masked gradients, apply them to the model’s neurons, and model them to forget the contribution of the specific samples. To implement this process, they introduced a mask gradient generator. The proposed mask gradient generator could adjust the volume of the mask gradients to avoid unlearning catastrophes based on the observation of the updated model.
@@
Sommer et al. (Sommer et al., 2022) introduced Athena, which leverages the property of backdoors, to verify the effectiveness of unlearning. Athena effectively certifies whether the data is deleted from an unlearning method with high confidence. Thus, it provides a basis for quantitatively inferring unlearning. In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is deleted.
7. Privacy Threats on Machine Unlearning

### v2

@@ 4.2.3. Bayesian-based Unlearning
Different from certified data removal that unlearns samples by subtracting corresponding Hessian matrix estimation from trained models, Bayesian-based unlearning tries to unlearn an approximate posterior as the model is trained by employing the remaining dataset. The exact Bayesian unlearning posterior can be derived from the Bayesian rule as $p(\theta|D_{r})=p(\theta|D)\ p(D_{e}|D_{r})/p(D_{e}|\theta)$, where $\theta$ is the posterior (i.e., model parameters). The erased dataset and the remaining dataset are two independent subsets of the full training dataset. If the model parameters $\theta$ are discrete-valued, $p(\theta|D_{r})$ can be directly obtained from the Bayesian rule (Nguyen et al., 2020). Additionally, employing a conjugate prior simplifies the unlearning process.
@@
Nevertheless, it is challenging to get the exact posterior in practice, not to mention the unlearning posterior. In (Nguyen et al., 2020), Nguyen et al. tried doing likewise at the beginning. They defined the loss function using the KLD between the approximate predictive distribution $q_{u}(y|D_{r})$ and the exact predictive distribution $p(y|D_{r})$. They bounded this loss function by the KLD between posterior beliefs $q_{u}(\theta|D_{r})$ and $p(\theta|D_{r})$ and further proposed evidence upper bound (EUBO) as the loss function to unlearn the approximate unlearning posterior. To avoid the overestimation of using KL divergence to optimize the posterior, they introduced an adjusted likelihood to control the unlearning extent.
@@
In (Nguyen et al., 2022a), the authors also studied the problem of ”unlearning” particular erased subset samples from a trained model with better efficiency than retraining a new model from scratch. Toward this purpose, Nguyen et al. (Nguyen et al., 2022a) proposed an MCMC-based machine unlearning method deriving from the Bayesian rule. They experimentally proved that MCMC-based unlearning could effectively and efficiently unlearn the erased subsets of the whole training dataset from a prepared model.
@@
Fu et al. (Fu et al., 2022) converted the MCMC unlearning problem into an explicit optimization problem. Then they proposed $\epsilon-$knowledge removal, which was a little similar to certified removal methods, but they defined that KLD between unlearned and retrained models must be less than $\epsilon$. To quantify the explicit $\epsilon-$knowledge removal, they proposed a knowledge removal estimator to assess the difference between the original and unlearned distributions. As they indicated, though their algorithm cannot wholly remove the learned knowledge from the already trained distribution, their method can still help the unlearned model approach its local minimum.
@@
Figure 7. (a) Regular data unlearning, which only unlearns the data sample. (b) Graph unlearning. Since the graph data contains both node features and connective edge information, we may need to unlearn edge information or both edges and nodes in graph unlearning.
4.2.4. Graph Unlearning
We introduce graph unlearning as a representative kind of irregular data unlearning. In (Chien et al., 2022a; Chen et al., 2022b; Cong and Mahdavi, 2022; Zhang et al., 2024), researchers extend regular data machine unlearning to a graph data scenario. Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig. 7. Therefore, Chien et al. (Chien et al., 2022a) proposed node unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC). Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees. They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
@@ 4.2.4. Graph Unlearning
Chen et al. (Chen et al., 2022b) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation. Therefore, they proposed a method called GraphEraser to implement unlearning tailored to graph data. Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs. Then, they trained the constituent graph models on these sub-graphs and ensembled them for the final prediction task. To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
@@
In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN. To remove the knowledge of a specified node, they design a projection-based unlearning approach, PROJECTOR, that projects the weights of the pre-trained model onto a subspace irrelevant to the deleted node features. PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
5. Federated Unlearning
We have introduced popular centralized machine unlearning in the former section. Their main purpose is to reduce the calculation and storage costs and the accuracy degradation caused by unlearning. Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in federated learning (FL) scenarios. In this section, we will introduce the literature about federated unlearning.
FL was initially introduced to protect the privacy of participating clients during the machine learning training process in distributed settings. All participants will only upload their locally trained model parameters instead of their sensitive local data to the FL server during model training processes. Therefore, in a federated learning scenario, limited access to the dataset will become a unique challenge when implementing unlearning.
Since the local data cannot be uploaded to the federated learning (FL) server side, most federated unlearning methods try to erase a certain client’s contribution from the trained model by storing and estimating the contribution of uploaded parameters. In this situation, they can implement federated unlearning without interacting with the client, shown as the server-side federated unlearning in Fig. 8 (a). The two representative methods are (Liu et al., 2021; Wu et al., 2022b). Liu et al. (Liu et al., 2021) proposed “FedEraser” to sanitize the impact of a FL client on the global FL model. In particular, during FL training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates. Based on these operations, they reconstructed the unlearned FL model instead of retraining a new model from scratch. However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data. It is unsuitable for a client who wants to unlearn a small piece of his data. Study (Wu et al., 2022b) tried to erase a client’s influence from the FL model by removing the historical updates from the global model. They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
Figure 8. Comparison between (a) server-side federated unlearning and (b) client-side federated unlearning
Wang et al. (Wang et al., 2022a) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL. The biggest challenge is that the local data utilized for the FL global model training cannot be accessed globally. Therefore, they explored the inner influence of each channel and observed that various channels have distinct impacts on different categories. They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model. A method called TF-IDF was introduced to quantize the class discrimination of channels. Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them. Evaluations of CIFAR10 and CIFAR100 demonstrate the unlearning effect of their method.
Different from unlearning a whole client’s influence and unlearning a class, Liu et al. (Liu et al., 2022a; Wang et al., 2023c) investigated how to unlearn data samples in FL, shown as the client-side federated unlearning in Fig. 8 (b). In (Liu et al., 2022a), they first defined a federated unlearning problem and proposed a fast retraining method to withdraw the influence of data from the FL model. Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods and the first-order Taylor approximate method (Liu et al., 2022a). They utilized the practical Fisher Information Matrix to model the Hessian matrix at a low cost. When unlearning, they subtract the estimation matrix from gradients with a bound. In (Wang et al., 2023c), they implement federated unlearning based on Bayesian inference, and they propose a parameters self-sharing method to reduce the model utility degradation.
@@ 6. Unlearning Evaluation and Verification
Table 3. Evaluation and Verification Metrics
Evaluation Metrics Description Usage Literature
Accuracy Model accuracy on erased datasets and remaining datasets To evaluate the predictive accuracy of the unlearned model (Bourtoule et al., 2021; Guo et al., 2020), …
@@
KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(\mathcal{A}(D_{r})||\mathcal{U}(D_{e},\mathcal{A}(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
JS-Divergence The distance between the predictions of retrained and unlearned model: $JS(\mathcal{A},\mathcal{U})=0.5\cdot KL(\mathcal{A}||Q)+0.5\cdot KL(\mathcal{U}||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
Membership inference Recall (##\##detected objects /#absent#/\#/ #erased objects) To verify if the erased sample is unlearned by the model (Graves et al., 2021)
@@
MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries To verify if the model unlearns the backdoored data (Hu et al., 2022)
Forgetting rate (FR) $FR=\frac{AF-BF}{BT}$, where AF, BF and BT are defined below To measure the rate of samples that are modified from member to non-member after unlearning (Ma et al., 2022)
$tr(I(\theta;D))$ is the trace of $I(\theta;D)$, and $I(\theta;D)$ is the Fisher Information matrix (Becker and Liebig, 2022).
6.1. The Unlearning Verification and Auditing Methods
@@
Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in machine unlearning. In Section 3, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics. The common metrics also include accuracy in assessing the performance of the unlearned model and running time in evaluating the unlearning efficiency. We have listed recent evaluation and verification metrics in Table 3, where the popular experimental datasets are furthermore presented. Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here. In this section, we introduce some new evaluation metrics that recent studies tailored to unlearning. Most of them are based on privacy leakage.
@@
In (Liu and Tsaftaris, 2020), Liu and Tsaftaris indicated that the data auditing problem in machine unlearning is becoming popular. They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed. Their challenge was that the method would fail when the calibration dataset was out of high quality, or the query samples were similar to the training dataset. The EMA (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model. EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools. In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset. EMA aggregates a final auditing score in the second step by ensembling many membership metrics and statistical tools. They verified whether a trained model forgets query data by the final auditing score. They conducted the experiments using standard datasets, including MNIST, SVHN, and the Chest X-ray dataset.
@@
Inspired by backdoor attacks in ML, Hu et al. (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB). MIB leverages the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong classes. The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if the model has been trained using her dataset. MIB evaluates the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model. Sequentially, they observed the results between the targeted model and clean models to infer whether the model was trained based on the backdoored samples.
@@
A similar membership-inference-based method was proposed in (Ma et al., 2022). Ma et al. (Ma et al., 2022) verified the effect of unlearning methods by the proposed forgetting rate (FR) metric. The evaluation metrics were defined using the observation of membership inference. Suppose $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$. In their FR definition, $BF$ and $AF$ are the samples in $D_{e}$ that are predicted as false by a membership inference attack before and after machine unlearning operations. $BT$ is the sample size in $D_{e}$, which is indicated to be correct by a membership inference attack before machine unlearning. According to the definition, $FR$ presents an instinctive evaluation of how many data points are altered from member to non-member by unlearning. If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective. By contrast, the unlearning will be meaningless. They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets. Forsaken’s main idea is to generate masked gradients continuously, apply them to the model’s neurons, and model them to forget the contribution of the specific samples. To implement this process, they introduced a mask gradient generator. The proposed mask gradient generator could adjust the volume to prevent unlearning catastrophes by observing the updated model.
@@
Sommer et al. (Sommer et al., 2022) introduced “Athena”, which leverages the property of backdooring techniques to verify the effectiveness of unlearning. Athena effectively and confidently certifies whether the data is deleted from an unlearning method. Thus, it provides a basis for quantitatively inferring unlearning. In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is unlearned.
Table 4. The Employed Datasets in Machine Unlearning
Data Type Name of Datasets Feature Dimension #. Smaples Task Type Employed by
Image MNIST (LeCun et al., 1998) $28\times 28\times 1$ $70,000$ Classification (Gupta et al., 2021; Wang et al., 2023a), …
CIFAR10 $32\times 32\times 3$ $60,000$ Classification (Wang et al., 2023a; Tarun et al., 2023), …
CIFAR100 $32\times 32\times 3$ $60,000$ Classification (Tarun et al., 2023; Cha et al., 2024), …
SVHN $32\times 32\times 3$ $99,289$ Classification (Bourtoule et al., 2021; Chundawat et al., 2023), …
ImageNet (Deng et al., 2009) $224\times 224\times 3$ $1,281,167$ Classification (Cha et al., 2024; Tarun et al., 2023), …
GTSRB (Stallkamp et al., 2011) $32\times 32\times 3$ $51,839$ Classification (Wu et al., 2022b; Wei et al., 2023), …
Market-1501 (Zheng et al., 2015) $128\times 64\times 3$ $32,668$ Person re-identification (Mehta et al., 2022)
Tabular Adult $14$ $48,842$ Classification (Chen et al., 2021; Warnecke et al., 2024), …
Credit info $30$ $284,807$ Classification and anomaly detection (Warnecke et al., 2024; Du et al., 2019), …
Covtype $54$ $581,012$ Classification (Mahadevan and Mathioudakis, 2021; Wu et al., 2023c), …
HIGGS (Baldi et al., 2014) $28$ $11,000,000$ Binary classification (Mahadevan and Mathioudakis, 2021, 2022), …
YELP2018 $5$ $1,561,406$ Recommendation (Chen et al., 2022a)
Movielens-1m $4$ $1,000,209$ Recommendation (Chen et al., 2022a)
Movielens-10m $4$ $10,000,054$ Recommendation (Chen et al., 2022a)
Text AG News (Zhang et al., 2015) $3$ $127,600$ Text classification (Wu et al., 2020)
RCV1 (Glewis et al., 2004) $3$ $804,414$ Text classification (Wu et al., 2020)
Graph Amazon Photo (Shchur et al., 2018) Features per Node: $745$ Nodes: $7,650$ Edges: $119,081$ Node classification Link prediction (Chien et al., 2022a, b), …
Cora (Sen et al., 2008) Features per Node: $1,433$ Nodes: $2,708$ Edges: $5,429$ Node classification Link prediction (Chien et al., 2022b; Chen et al., 2022b), …
Citseer (Yang et al., 2016) Features per Node: $3,703$ Nodes: $3,327$ Edges: $4,732$ Node classification Link prediction (Chen et al., 2022b)
Pubmed (Fey and Lenssen, 2019) Features per Node: $500$ Nodes: $19,717$ Edges: $44,338$ Node classification Link prediction (Chen et al., 2022b; Chien et al., 2022a), …
ogbn-arxiv (Hu et al., 2020) Features per Node: $128$ Nodes: $169,343$ Edges: $1,166,243$ Node classification Link prediction (Chien et al., 2022b, a), …
Computers (McAuley et al., 2015) Features per Node: $767$ Nodes: $13,752$ Edges: $245,861$ Node classification Link prediction (Chen et al., 2022b; Chien et al., 2022a), …
CS (Shchur et al., 2018) Features per Node: $767$ Nodes: $18,333$ Edges: $327,476$ Node classification Link prediction (Chen et al., 2022b)
Physics (Shchur et al., 2018) Features per Node: Text Nodes: $27,770$ Edges: $352,807$ Node classification Link prediction (Chen et al., 2022b)
6.2. The Employed Datasets
We collect the commonly employed datasets in machine unlearning studies and present the detail introduction of them in Table 4. There are four main types of data: Image, Tabular, Text and Graph. Most of the unlearning studies use image datasets and train classification models based on these image datasets. For tabular datasets, most of them are used in recommendation systems. The unlearning studies that investigate how to unlearn a recommendation model will use these tabular datasets. Graph data is employed for node classification and link prediction tasks, which is usually used in graph unlearning studies. For convenience to find the related studies, we link the corresponding unlearning studies at the last column in Table 4.
7. Privacy and Security Threats on Machine Unlearning